A person traveling 4 hr by plane and 25 hr by ship covers 1580 miles. If the speed of the plane had been one-half of the actual speed and the speed of the ship had been one-fourth greater, the person would have traveled only 1315 miles in the same length of time. Find the speeds of the plane and the ship..
Having a problem with this one ......thank you..

Answer by stanbon(75887)

A person traveling 4 hr by plane and 25 hr by ship covers 1580 miles. If the speed of the plane had been one-half of the actual speed and the speed of the ship had been one-fourth greater, the person would have traveled only 1315 miles in the same length of time. Find the speeds of the plane and the ship..
--------------------------
Original Plane DATA:
Rate = x mph ; Time = 4 hr. : Distance = rt = 4x miles
--------------
Original Ship DATA:
Distance = 1580-4x miles ; Time = 25 hrs. ; Rate = d/t = (1580-4x)/25 mph
------------------
New Plane DATA:
Rate = x/2 mph ; Time = 4 hr ; Distance = rt = 2x miles
------
New Ship DATA:
Rate = (3/2)(1580-4x)/25 mph ; Time = 25 hrs. ; Dist. = rt = (3/2)(1580-4x) mi.
---------------------
EQUATION:
(3/2)(1580-4x) = 1315
1580-4x = 876.33333333
4x = 703.333333333
x = 175.8333333...(plane speed)
-------------
Ship speed = (1580-4x)/25 mph
(1580-4*175.833333333)/25) = 35.07 mph (ship speed)
========================
Cheers,
Stan H.